the probability that a student of a school owns a car is 1/5, three students were sampled, Find the probability that only one of them owns a car?
1
Expert's answer
2012-12-06T11:12:45-0500
At first we pick one studentamong these three who has a car (3 possible ways); the probability in any of
these cases is
P=1/5(car)*4/5(no car)*4/5(no car)=0.128. Total probability is Number_of _possible_ways*probability_of such_event=3*0.128=0.384